OpenStudy (anonymous):

given x^2 + y^2 = 25, find d^2y/dx^2

OpenStudy (anonymous):

i can find the first derivative (-x/y)

OpenStudy (anonymous):

2x + 2y (dy/dx) = 0 So dy/dx = -x/y Take derivative again: 2 + (2y)(d^2y/dx^2) + 2 (dy/dx) = 0 d^2y/dx^2 = (-2(-x/y) -2) / 2y = (x-y)/y^2